Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-12-02

Observers: Miika Pursiainen &  Noel  Castro Segura
Support: Claudia Gutiérrez & Tomás Müller Bravo
TAT: Joe Anderson  & Janet Chen


Calibrations:
EFOSC:
biases
dome flat fields
arcs (evening and morning)
UVRi Sky Flats (morning), B band had to be aborted

Weather:
Start:      Thin clouds, seeing 1.0-1.5 and varies a lot, windy, but telescope remains open
Middle:  Thin clouds, seeing 0.7-1.3
End:      Thin clouds, seeing 0.9-1.4

Problems during the night:
- There has been problems with the secondary mirror causing the image focusing to be take time. Sometimes the focus is not properly made causing the images to have a bit of coma.

Instrument - EFOSC

UT            Target               Type                Setup         airmass             Seeing              notes
-----------------------------------------------------------------------------------------
                 Image Focusing
00:30        SN2018jao       class                gr13            1.5                     1.2                     Ia few days before peak
00:45        SN2018fit         follow              gr13            1.5-1.8               1.3-1.5
02:30        VMA2               STD                  all               1.4                      1.0
02:55        AT2018fyk        follow              gr11            1.5-2.5               1.1-1.02      
                 Image Analysis
04:40        SN2018rw        follow              Vri               1.05-1.02           1.2                     Telescope had to rotate 360 degrees during i band exposures (too close to culmination)
05:55        L745a               STD                  all               1.1                      0.9
06:25        SN2018hle       follow              gr13            1.4                      1.3
06:55        AT2018hyz       follow              gr11+16      1.7                      1.4
07:55        AT2018jdq       class                gr13            1.6                      1.2                    Transient not seen in the acquisition image
08:05        L745a               STD                  all                1.1                      1.1
Comments